X-2y=6
y=2X-18
Solve the system by the substitution method.
Please help me solve.
Thanks, angela

Found 2 solutions by sofiyac, Mr.Math:
Answer by sofiyac(983)

plug in the second one into the first one for y
so you get
x-2(2x-18)=6
x-4x+36=6
-3x+36=6
-3x=-33
x=11
now plug this back into original second equation
y=2x-18=2*11-18=22-18=4
solution (11,4)

Answer by Mr.Math(29)

X-2y=6
y=2X-18
x=2(2x-18)+6
x=4x-36+6
3x=30
x=10 ; y=20-18=2